Use the information in the table below to answer the questions concerning Company C’s current credit policy.

net 40

Sales

$9,000,000

Lag factor

3

Cost of capital

8%

Bad debt, % sales

5%

Cost of goods sold, % sales

79%

G&A, % sales

9%

1.         What is the present value of the firm’s daily net sales?

2.         What is the firm’s daily cost of good sold?

3.         What is the firm’s DSO (Net + lag factor)?

4.         What is the present value of the firm’s daily G&A expense?

5.         What is the net present value of the firm’s current policy?
Answer #1

Assuming the above figure is given on per annum (360 days) following is done:

1) PV of the firm's daily net Sales = Daily net sales * PV factor for the net credit period = ($9,000,000*0.95/360) * 1/(1+0.08*43/360) = $23,525.20 (approx).

2) Firm's daily cost of goods sold = Firm's daily sales * COGS % = $9,000,000/360 * 79% = $19,750.

3) Firm's Days Sales Outstanding (Net + lag factor) = 40+3 = 43 days.

4) PV of the firm's daily G&A expenses = Daily G & A expenses * PV factor for the period = $9,000,000/360 * 9% * 1/(1+0.08) = $2,083.33.